# Novel scaling laws to derive spatially resolved flare and CME parameters from sun-as-a-star observables

### v1

@@
Coronal mass ejections (CMEs) are often associated with X-ray (SXR) flares powered by magnetic reconnection in the low-corona, while the CME shocks in the upper corona and interplanetary (IP) space accelerate electrons often producing the type-II radio bursts. The CME and the reconnection event are part of the same energy release process as highlighted by the correlation between reconnection flux ($\mathrm{\phi_{rec}}$) that quantifies the strength of the released magnetic free energy during SXR flare, and the CME kinetic energy that drives the IP shocks leading to type-II bursts. Unlike the sun, these physical parameters cannot be directly inferred in stellar observations. Hence, scaling laws between unresolved sun-as-a-star observables, namely SXR luminosity ($\mathrm{L_{X}}$) and type-II luminosity ($\mathrm{L_{R}}$), and the physical properties of the associated dynamical events are crucial. Such scaling laws also provide insights into the interconnections between the particle acceleration processes across low-corona to IP space during solar-stellar ‘flare- CME- type-II’ events. Using long-term solar data in SXR to radio waveband, we derive a scaling law between two novel power metrics for the flare and CME-associated processes. The metrics of ‘flare power’ ($\mathrm{P_{flare}}$=$\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and ‘CME power’ ($\mathrm{P_{CME}}$=$\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$), where $\mathrm{V_{CME}}$ is the CME speed, scale as $\mathrm{P_{flare}}$$\propto$ $\mathrm{P_{CME}}$0.76±0.04. Besides, $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ show power-law trends with $\mathrm{P_{CME}}$ with indices of 1.12$\pm$0.05 and 0.61$\pm$0.05 respectively. These power-laws help infer the spatially resolved physical parameters, $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$, from disk-averaged observables, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ during solar-stellar ‘flare- CME- type-II’ events.
@@ 1 Introduction
Coronal mass ejections (CMEs) are a major driver of space weather events in solar and stellar environments (see, Howard et al., 2023, for an overview). A CME can produce shocks across corona out to interplanetary (IP) space, accelerating electrons which often produce meter-kilometer (m-km) band type-II bursts (Wild & McCready, 1950; Wild, 1970; McLean & Labrum, 1985; Gopalswamy, 2011; Miteva et al., 2017; Alvarado-Gómez et al., 2022). Depending on the heights of shock formation, the type-II bursts may have a start and end frequency anywhere within the m-km range. The CMEs associated with the type-II bursts are generally much faster and wider than an average CME, with 78% of them associated with major solar energetic particle (SEP) events (Gopalswamy et al., 2005; Kumari et al., 2023). The type-II associated CMEs are also accompanied by relatively intense soft X-ray (SXR) flares (Gopalswamy, 2006). Hence, the type-II bursts are highly sought after in active stars as a CME shock signature (e.g. Osten & Bastian, 2008; Crosley & Osten, 2018; Villadsen et al., 2014).
@@
When the type-II burst conveys the strength of particle acceleration caused by the CME shocks, the associated SXR flare is produced by the heating of the post-reconnnection magnetic field structures in the low-corona. Hence the peak SXR luminosity ($\mathrm{L_{X}}$) correlates well with the reconnection flux ($\mathrm{\phi_{rec}}$), which is a proxy to the strength of the reconnection-driven currents that drive the local heating (Kazachenko et al., 2017; Sindhuja & Gopalswamy, 2020). Gopalswamy et al. (2018a) showed that $\mathrm{\phi_{rec}}$ correlates well with CME kinetic energy, revealing a link between the low-coronal and IP space impacts of the energy release event, of which the CME and the flare are part of (hereafter, ‘flare- CME- type-II’ event). However, there exist no scaling laws involving $\mathrm{L_{X}}$, type-II luminosity ($\mathrm{L_{R}}$), and the physical properties of the flare and the CME, that could reveal any links between the mechanisms and drivers of particle acceleration across corona to IP space. Besides, since the flare and type-II burst outshine the quiet sun during a ‘flare- CME- type-II’ event, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ are sun-as-a-star observables. So, the scaling laws of the aforementioned kind are relevant for stellar CME studies, that lack spatially resolved observations. In this context, the Güedel-Benz relationship (GBR; Gudel et al., 1993) that connects $\mathrm{L_{X}}$ and microwave (5 - 8 GHz) luminosity by a power-law is noteworthy. GBR revealed the link between the population of flare-accelerated electrons propagating towards the sun causing the microwave and SXR flare, and it provides a common framework to explore the solar and stellar flares (Guedel & Benz, 1993; Guedel et al., 1995; Airapetian & Holman, 1998). This work will explore analogous correlations between $\mathrm{L_{X}}$, $\mathrm{L_{R}}$, and the physical properties of the ‘flare- CME- type-II’ events.
@@ 2 Data and Methodology
We use the calibrated decameter-hectrometric (DH) radio dynamic spectra from Radio and Plasma Wave Investigation (WAVES) instruments onboard Wind, STEREO A and STEREO B spacecraft. The DH type-IIs were chosen for this study over the metric bursts for multiple reasons. Firstly, the DH type-II bursts are produced at regions of CME-driven interplanetary shocks caused by relatively stronger flares and faster CMEs than their metric counterparts (Gopalswamy, 2011; Miteva et al., 2017). Also, DH bursts are more closely associated with SEPs and sustained gamma ray emission than the type-IIs confined within the metric band, making these bursts more interesting from a space weather perspective (Gopalswamy, 2006; Miteva et al., 2017; Gopalswamy et al., 2018b). The database of the calibrated DH dynamic spectra from each spacecraft gathered over multiple solar cycles forms a uniformly calibrated long-term dataset. Besides, simultaneous STEREO and Wind observations form a unique multi-vantage point radio database.
@@
Figure 1: (a) $\mathrm{L_{R}}$ histogram for DH type-II event lists and the full catalog. Sample sizes are given in square brackets. The distributions look similar across event lists with a median flux density of $10^{4.5}$ SFU. (b) Properties of List 2 events. $\mathrm{L_{X}}$ is not well correlated with $\mathrm{L_{R}}$.
@@
The events reported in the DH type-II catalog (Gopalswamy et al., 2019) between Nov 2006 and July 2023 form our initial sample. For each event, the data from all spacecraft were examined and the best-recorded event dynamic spectrum (DS) was chosen for further analysis. The peak flux of the burst within 3 - 7 MHz range was used to compute $\mathrm{L_{R}}$. Within 3 - 7 MHz the major burst types include type-II and type-III bursts. The long-duration broadband type-IV bursts are mostly confined above 7 MHz (Mohan et al., 2024a). Type-IIIs have very high frequency-drift rates than type-IIs making it easier to disentangle the emission from near-simultaneous bursts in the DS. So, 3 - 7 MHz range is an ideal choice to ensure robust flux estimation in most cases with minimal impact from co-temporal bursts. $\mathrm{L_{X}}$ and CME parameters were obtained by cross-matching the radio event list with the publicly available catalogs maintained by the Coordinated Data Analysis Workshops (CDAW) group [^1] https://cdaw.gsfc.nasa.gov/. The resultant catalog (hereafter ‘full catalog’) has 163 bursts with 3- 7  MHz band $\mathrm{L_{R}}$ estimates. Of these, 118 events (List 1) have good estimation of $\mathrm{L_{X}}$ and CME properties, especially its speed ($\mathrm{V_{CME}}$). From List 1, the type-II bursts that are well isolated in the DS, without potential contamination from other emission features were chosen. $\mathrm{\phi_{rec}}$ was estimated for all possible events using the post eruption arcade (PEA) method described in Gopalswamy et al. (2017). All these resulted in a list of 25 ‘flare- CME- type-II’ events (List 2) with reliable estimates for $\mathrm{L_{X}}$,$\mathrm{L_{R}}$,$\mathrm{V_{CME}}$, and $\mathrm{\phi_{rec}}$, which will be the focus of our analysis. The full catalog, event lists, the calibrated DS across the entire WAVES band and within 3 - 7 MHz, and a detailed description of the tables are made available online [^2] https://cdaw.gsfc.nasa.gov/CME_list/radio/multimission_type2/.
@@ 3 Results
Figure 2: (a) Spearman Correlation matrix. Each cell provides the Spearmann cross-correlation coefficient (CR,C) between the respective row (R) and column (C) parameter. Cell color represents the p-value. The CR,C, error and p-value are found using a Bootstrapping analysis. The ‘very significant’ (p$<$0.015), ‘significant’ (p$<$0.05) and ‘insignificant’ (p$>$0.05) CR,C values are shown in green, white and red texts respectively. The CR,C between the novel composite metrics and the respective dependent parameters are masked and marked by ‘–’. (b) $\mathrm{P_{flare}}$- $\mathrm{P_{CME}}$scaling law. (c-d) Relations connecting $\mathrm{\phi_{rec}}$ and $\mathrm{V_{CME}}$ to $\mathrm{P_{CME}}$. Shaded regions represent the fitting uncertainty intervals.
@@
Figure 1a shows the histogram of $\mathrm{L_{R}}$ for the full catalog and the sub-lists. The median flux density is $\sim 3.5\times 10^{4}$ SFU for all lists. Figure 1b shows the flare and CME properties of List 2 events. There is no significant correlation between $\mathrm{L_{R}}$ and $\mathrm{L_{X}}$. However, certain other parameter pairs appear to be correlated, which will be discussed in detail. Two novel metrics of ‘flare power’ ($\mathrm{P_{flare}}$= $\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and ‘CME power’ ($\mathrm{P_{CME}}$= $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$) are defined by combining $\mathrm{\phi_{rec}}$ and $\mathrm{V_{CME}}$, with $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ respectively. Spearman cross-correlation analysis was carried out between each of the parameters considered, namely $\mathrm{L_{X}}$,$\mathrm{L_{R}}$,$\mathrm{\phi_{rec}}$,$\mathrm{V_{CME}}$,$\mathrm{P_{flare}}$ and $\mathrm{P_{CME}}$. Given the small sample size, a bootstrap technique was employed to obtain robust correlation coefficients and p-values. Of the 25 events, 18 ($\sim$70%) were randomly selected iteratively while computing cross-correlations, CR,C, between each parameter pair (R,C) over 10000 rounds. The median of the 10000 CR,C and p vales for each pair was noted. The errors on each median CR,C is the median absolute deviation. Fig. 2a shows the cross-correlation matrix with each cell providing the median CR,C between the respective row (R) and column (C) parameters. The cell color denote the median p-value, and the text color of CR,C denotes the statistical significance. ||||CR,C|||| values with p$<$0.015 are considered very significant (green text), while 0.015 $<$p$<$0.05 are significant (white text) and those with p$>$0.05 are insignificant (red text). The CR,C for the tuples ($\mathrm{P_{flare}}$,$\mathrm{L_{X}}$), ($\mathrm{P_{flare}}$,$\mathrm{\phi_{rec}}$), ($\mathrm{P_{CME}}$,$\mathrm{L_{R}}$) and ($\mathrm{P_{CME}}$,$\mathrm{V_{CME}}$) are ignored (marked by ‘–’) since the row parameter is a function of the column parameter by definition.
@@
The parameter tuples with significant correlation include ($\mathrm{\phi_{rec}}$,$\mathrm{L_{X}}$), ($\mathrm{P_{CME}}$,$\mathrm{P_{flare}}$), ($\mathrm{P_{CME}}$,$\mathrm{L_{X}}$) and ($\mathrm{P_{CME}}$,$\mathrm{\phi_{rec}}$). Our $\mathrm{C_{L_{X},\phi_{rec}}}$ matches the recent result by Kazachenko (2023). However, both $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ are related to the low-coronal flare, while we aim to identify links between the powering of the low-coronal flare and the interplanetary type-II burst. This makes the other significantly correlated tuples more interesting.
3.1 Flare - CME - type-II scaling laws
@@
Of all possible parameter pairs where one is purely related to the SXR flare and the other to radio burst physics, the engineered novel power metrics ($\mathrm{P_{flare}}$,$\mathrm{P_{CME}}$) shows the highest correlation. We derive a new robust scaling law between the powers in the flare- and CME-driven activity,
@@
Interestingly, the power-law index coincides with that of GBR ($\mathrm{L_{X}}$ $\propto$ $\mathrm{L_{R}}$0.73; Guedel et al., 1995) within estimation errors. GBR computes $\mathrm{L_{R}}$ in the 5 - 8 GHz range. The other significant strongly correlated tuples also provide useful scaling laws;
@@
Note that $\mathrm{V_{CME}}$ and $\mathrm{L_{R}}$ that make up $\mathrm{P_{CME}}$, have insignificant or no correlation with $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$, making the strong $\mathrm{P_{CME}}$- $\mathrm{P_{flare}}$ correlation robust. The power-law indices are constrained fairly well. More data from the ongoing solar cycle will help better constrain both the fit parameters in the future.

### v2

@@
Coronal mass ejections (CMEs) are often associated with X-ray (SXR) flares powered by magnetic reconnection in the low corona, while the CME shocks in the upper corona and interplanetary (IP) space accelerate electrons often producing the type II radio bursts. The CME and the reconnection event are part of the same energy release process as highlighted by the correlation between reconnection flux ($\mathrm{\phi_{rec}}$) that quantifies the strength of the released magnetic free energy during the SXR flare, and the CME kinetic energy that drives the IP shocks leading to type II bursts. Unlike the Sun, these physical parameters cannot be directly inferred in stellar observations. Hence, scaling laws between unresolved sun-as-a-star observables, namely SXR luminosity ($\mathrm{L_{X}}$) and type II luminosity ($\mathrm{L_{R}}$), and the physical properties of the associated dynamical events are crucial. Such scaling laws also provide insights into the interconnections between the particle acceleration processes across low-corona to IP space during solar-stellar ”flare-CME-type II” events. Using long-term solar data in the SXR to radio waveband, we derived a scaling law between two novel power metrics for the flare and CME-associated processes. The metrics of ”flare power” ($\mathrm{P_{flare}}$=$\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and ”CME power” ($\mathrm{P_{CME}}$=$\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$), where $\mathrm{V_{CME}}$ is the CME speed, scale as $\mathrm{P_{flare}}$$\propto$ $\mathrm{P_{CME}}$0.76±0.04. In addition, $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ show power-law trends with $\mathrm{P_{CME}}$ with indices of 1.12$\pm$0.05 and 0.61$\pm$0.05, respectively. These power laws help infer the spatially resolved physical parameters, $\mathrm{V_{CME}}$ and $\mathrm{\phi_{rec}}$, from disk-averaged observables, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$  during solar-stellar flare-CME-type II events.
@@ 1 Introduction
Coronal mass ejections (CMEs) are a major driver of space weather events in solar and stellar environments (see, Howard et al., 2023, for an overview). A CME can produce shocks across corona out to interplanetary (IP) space, accelerating electrons which often produce meter-kilometer (m-km) band type II bursts (Wild & McCready, 1950; Wild, 1970; McLean & Labrum, 1985; Gopalswamy, 2011; Miteva et al., 2017; Alvarado-Gómez et al., 2022). Depending on the heights of shock formation, the type II bursts may have a start and end frequency anywhere within the m-km range. The CMEs associated with the type II bursts are generally much faster and wider than an average CME, with 78% of them associated with major solar energetic particle (SEP) events (Gopalswamy et al., 2005; Kumari et al., 2023). The type II associated CMEs are also accompanied by relatively intense soft X-ray (SXR) flares (Gopalswamy, 2006). Hence, the type II bursts are highly sought after in active stars as a CME shock signature (e.g., Osten & Bastian, 2008; Crosley & Osten, 2018; Villadsen et al., 2014).
@@
When the type II burst conveys the strength of particle acceleration caused by the CME shocks, the associated SXR flare is produced by the heating of the post-reconnnection magnetic field structures in the low corona. Hence the peak SXR luminosity ($\mathrm{L_{X}}$) correlates well with the reconnection flux ($\mathrm{\phi_{rec}}$), which is a proxy for the strength of the reconnection-driven currents that drive the local heating (Kazachenko et al., 2017; Sindhuja & Gopalswamy, 2020). Gopalswamy et al. (2018a) show that $\mathrm{\phi_{rec}}$ correlates well with CME kinetic energy, revealing a link between the low-coronal and IP space impacts of the energy release event, of which the CME and the flare are a part (hereafter, flare-CME-type II event). However, there exist no scaling laws involving $\mathrm{L_{X}}$, type II luminosity ($\mathrm{L_{R}}$), and the physical properties of the flare and the CME, which could reveal any links between the mechanisms and drivers of particle acceleration across corona to IP space. In addition, since the flare and type II burst outshine the quiet sun during a flare-CME-type II event, $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ are sun-as-a-star observables. So, the scaling laws of the aforementioned kind are relevant for stellar CME studies, which lack spatially resolved observations. In this context, the Güedel-Benz relationship (GBR; Gudel et al., 1993) that connects $\mathrm{L_{X}}$ and microwave (5 - 8 GHz) luminosity by a power law is noteworthy. GBR revealed the link between the population of flare-accelerated electrons propagating toward the Sun causing the microwave and SXR flare, and it provides a common framework to explore the solar and stellar flares (Guedel & Benz, 1993; Guedel et al., 1995; Airapetian & Holman, 1998). This work explores analogous correlations between $\mathrm{L_{X}}$, $\mathrm{L_{R}}$, and the physical properties of the flare-CME-type II events.
@@ 2 Data and methodology
We used the calibrated decameter-hectometric (DH) radio dynamic spectra from Radio and Plasma Wave Investigation (WAVES) instruments on board Wind, STEREO A, and STEREO B spacecraft. The DH type IIs were chosen for this study over the metric bursts for multiple reasons. Firstly, the DH type II bursts are produced at regions of CME-driven interplanetary shocks caused by relatively stronger flares and faster CMEs than their metric counterparts (Gopalswamy, 2011; Miteva et al., 2017). Also, DH bursts are more closely associated with SEPs and sustained $\gamma$-ray emission than the type IIs confined within the metric band, making these bursts more interesting from a space weather perspective (Gopalswamy, 2006; Miteva et al., 2017; Gopalswamy et al., 2018b). The database of the calibrated DH dynamic spectra from each spacecraft gathered over multiple solar cycles forms a uniformly calibrated long-term dataset. In addition, simultaneous STEREO and Wind observations form a unique multi-vantage point radio database.
@@
Figure 1: Statistical properties of flare-CME-type II events. (a) The $\mathrm{L_{R}}$ histogram for DH type II event lists and the full catalog. Sample sizes are given in square brackets. The distributions look similar across event lists with a median flux density of $10^{4.5}$ SFU. (b) Properties of List 2 events. The $\mathrm{L_{X}}$ is not well correlated with $\mathrm{L_{R}}$.
@@
The events reported in the DH type II catalog (Gopalswamy et al., 2019) between November 2006 and July 2023 constitute the initial sample. For each event, data from all spacecraft were examined and the best-recorded event dynamic spectrum (DS) was chosen for further analysis. The peak flux of the burst within the 3 - 7 MHz range was used to compute $\mathrm{L_{R}}$. Within 3 - 7 MHz the major burst types include type II and type III bursts. The long-duration broadband type-IV bursts are mostly confined above 7 MHz (Mohan et al., 2024a). Type IIIs have very high frequency-drift rates in comparison to type IIs, making it easier to disentangle the emission from near-simultaneous bursts in the DS. As a result, the 3 - 7 MHz range is an ideal choice to ensure robust flux estimation in most cases with minimal impact from co-temporal bursts. The $\mathrm{L_{X}}$ and CME parameters were obtained by cross-matching the radio event list with the publicly available catalogs maintained by the Coordinated Data Analysis Workshops (CDAW) group [^1] https://cdaw.gsfc.nasa.gov/. The resultant catalog (hereafter ”full catalog”) has 163 bursts with 3- 7  MHz band $\mathrm{L_{R}}$ estimates. Of these, 118 events (List 1) have a good estimation for $\mathrm{L_{X}}$ and CME properties, especially their speed ($\mathrm{V_{CME}}$). From List 1, the type II bursts that are well isolated in the DS, without potential contamination from other emission features, were chosen. The $\mathrm{\phi_{rec}}$ was estimated for all possible events using the post-eruption arcade (PEA) method described in Gopalswamy et al. (2017). All of these resulted in a list of 25 flare-CME-type II events (List 2) with reliable estimates for $\mathrm{L_{X}}$, $\mathrm{L_{R}}$, $\mathrm{V_{CME}}$, and $\mathrm{\phi_{rec}}$, which are the focus of our analysis. The full catalog, event lists, the calibrated DS across the entire WAVES band and within 3 - 7 MHz, and a detailed description of the tables are made available online [^2] https://cdaw.gsfc.nasa.gov/CME_list/radio/multimission_type2/ and via Zenodo [^3] https://doi.org/10.5281/zenodo.13896427.
@@ 3 Results
Figure 2: Cross-correlations and scaling laws between various properties of flare-CME-type II events. (a) Spearman correlation matrix. Each cell provides the Spearmann cross-correlation coefficient (CR,C) between the respective row (R) and column (C) parameters. The cell color represents the p value. The CR,C, error, and p-value were found using a Bootstrapping analysis. The ”very significant” (p$<$0.015), ”significant” (p$<$0.05), and ”insignificant” (p$>$0.05) CR,C values are shown in green, white, and red text, respectively. The CR,C between the novel composite metrics and the respective dependent parameters are masked and marked by ”–”. (b) The $\mathrm{P_{flare}}$- $\mathrm{P_{CME}}$ scaling law. (c-d) Relations connecting $\mathrm{\phi_{rec}}$ and $\mathrm{V_{CME}}$ to $\mathrm{P_{CME}}$. Shaded regions represent the fitting uncertainty intervals.
3.1 Statistics and cross-correlation analysis
@@
Figure 1a shows the histogram of $\mathrm{L_{R}}$ for the full catalog and the sub-lists. The median flux density is $\sim 3.5\times 10^{4}$ SFU for all lists. Figure 1b shows the flare and CME properties of List 2 events. There is no significant correlation between $\mathrm{L_{R}}$ and $\mathrm{L_{X}}$. However, certain other parameter pairs appear to be correlated, which are subsequently discussed in detail. Two novel metrics of flare power ($\mathrm{P_{flare}}$= $\sqrt{\mathrm{L_{X}\phi_{rec}}}$) and CME power ($\mathrm{P_{CME}}$= $\sqrt{\mathrm{L_{R}V^{2}_{CME}}}$) are defined by combining $\mathrm{\phi_{rec}}$ and $\mathrm{V_{CME}}$, with $\mathrm{L_{X}}$ and $\mathrm{L_{R}}$ , respectively. Spearman cross-correlation analysis was carried out between each of the parameters considered, namely $\mathrm{L_{X}}$, $\mathrm{L_{R}}$, $\mathrm{\phi_{rec}}$, $\mathrm{V_{CME}}$, $\mathrm{P_{flare}}$ and $\mathrm{P_{CME}}$. Given the small sample size, a bootstrap technique was employed to obtain robust correlation coefficients and p values. Of the 25 events, 18 ($\sim$70%) were randomly selected iteratively while computing cross correlations, CR,C, between each parameter pair (R,C) over 10000 rounds. The median of the 10000 CR,C and p values for each pair were noted. The errors on each median CR,C is the median absolute deviation. Fig. 2a shows the cross-correlation matrix with each cell providing the median CR,C between the respective row (R) and column (C) parameters. The cell color denotes the median p value, and the text color of CR,C denotes the statistical significance. We note that ||||CR,C|||| values with p$<$0.015 are considered very significant (green text), while 0.015 $<$p$<$0.05 are significant (white text), and those with p$>$0.05 are insignificant (red text). The CR,C for the tuples ($\mathrm{P_{flare}}$,$\mathrm{L_{X}}$), ($\mathrm{P_{flare}}$,$\mathrm{\phi_{rec}}$), ($\mathrm{P_{CME}}$,$\mathrm{L_{R}}$), and ($\mathrm{P_{CME}}$,$\mathrm{V_{CME}}$) have been ignored (marked by ”–”) since the row parameter is a function of the column parameter by definition.
@@
The parameter tuples with a significant correlation include ($\mathrm{\phi_{rec}}$,$\mathrm{L_{X}}$), ($\mathrm{P_{CME}}$,$\mathrm{P_{flare}}$), ($\mathrm{P_{CME}}$,$\mathrm{L_{X}}$), and ($\mathrm{P_{CME}}$,$\mathrm{\phi_{rec}}$). Our $\mathrm{C_{L_{X},\phi_{rec}}}$ matches the recent result by Kazachenko (2023). However, both $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$ are related to the low-coronal flare, while we aim to identify links between the powering of the low-coronal flare and the interplanetary type II burst. This makes the other significantly correlated tuples more interesting.
3.2 Flare-CME-type II scaling laws
@@
Of all possible parameter pairs where one is purely related to the SXR flare and the other to radio burst physics, the engineered novel power metrics shows the highest correlation. We derived a new robust scaling law between the powers in the flare- and CME-driven activity,
@@
Interestingly, the power-law index coincides with that of GBR ($\mathrm{L_{X}}$ $\propto$ $\mathrm{L_{R}}$0.73; Guedel et al., 1995) within estimation errors. GBR computes $\mathrm{L_{R}}$ in the 5 - 8 GHz range. The other significant, strongly correlated tuples also provide useful scaling laws;
@@
We note that $\mathrm{V_{CME}}$ and $\mathrm{L_{R}}$ , which make up $\mathrm{P_{CME}}$, have insignificant or no correlation with $\mathrm{L_{X}}$ and $\mathrm{\phi_{rec}}$, making the strong $\mathrm{P_{CME}}$- $\mathrm{P_{flare}}$ correlation robust. The power-law indices are constrained fairly well. More data from the ongoing solar cycle will help better constrain both the fit parameters in the future.